OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2. A. y = 1/2 x + 2 B. y = 1/2 x C. y = 1/2 x + 4

OpenStudy (anonymous):

y-y1=m(x-x1).So here y-2=1/2(x-4).Simplify to get y=1/2x

OpenStudy (anonymous):

A. 1/2 is slope 2 is y-int